Griffin's Goat Farm, Inc., has sales of $682,000, costs of $344,000, depreciation expense of $88,000, interest expense of $54,000, a tax rate of 25 percent, and paid out $39,000 in cash dividends.

What is the addition to retained earnings?
Answer #1

Calculation of addition to retained earnings:

Net income before tax= sales-cost-depreciation-interest

Net income before tax= 682000-344000-88000-54000= 196000

Net income after tax= 196000*(1-0.25)= 196000*0.75= 147000

Addition to retained earning= net income after tax- dividend

Addition to retained earning= 147000-39000= 108000